OpenStudy (selena_gomez):

Which statement best explains the value of 16 − (−3)? The additive inverse of −3 is −3, so 16 − (−3) = 19. The additive inverse of −3 is −3, so 16 − (−3) = 13. The additive inverse of −3 is +3, so 16 − (−3) = 19. The additive inverse of −3 is +3, so 16 − (−3) = 13.

OpenStudy (kittykatiehunt):

is one of those a 12 instead of a 13 in the answers

OpenStudy (kittykatiehunt):

if not then it is c

OpenStudy (mathstudent55):

What is the additive inverse of a number?

OpenStudy (mathstudent55):

Also, subtraction is adding the additive inverse. To subtract a - b, add the additive inverse of b to a, a - (-b).

OpenStudy (skullpatrol):

Replace a with 16 and b with –3 in the definition of subtraction: $$\huge a-b=a+(-b)$$ $$\huge 16-(-3)=16+(-(-3))$$ $$\huge 16-(-3)=16+(+3)$$ $$\huge 16-(-3)=16+3$$

OpenStudy (skullpatrol):

To subtract b, add the opposite of b.

OpenStudy (skullpatrol):

opposite of a number: Each of the numbers in a pair such as 6 and –6 or –2.5 and 2.5. Also called additive inverse. The opposite of a number b is written as –b. Caution: –b is not necessarily a negative number. For example, if b = –3, then $$\huge –b = –(–3) = +3 = 3$$